OpenStudy (anonymous):

Prove the following Identity: sec^2x-1/cotx = tan^3x Help would be very welcomed thank you.

hero (hero):

Begin with the LHS:\[\frac{\sec^2x - 1}{\cot x} \] Substitute \(\sec^2x - 1 = \tan^2x\) \[\frac{\tan^2x}{\cot x}\] Write \(\frac{\tan^2x}{\cot x}\) as \[\tan^2x \div \cot x\]Which becomes \[\tan^2x \times \tan x\]